OpenStudy (anonymous):

A positive integer is 5 more than 23 times another. Their product is 6732. Find the two integers.

OpenStudy (anonymous):

x=34. What is the other number?

OpenStudy (misty1212):

lol another one!!

OpenStudy (misty1212):

if one is \(34\) (i didn't check, just assuming) then the other is \[23\times 34+5\]

OpenStudy (anonymous):

I only have about 10 more! :(

OpenStudy (misty1212):

lolololol

OpenStudy (misty1212):

oops it is wrong!!

OpenStudy (misty1212):

the answer is not 34

OpenStudy (anonymous):

how is it not?

OpenStudy (misty1212):

\[x(23x+5)=6732\] right?

OpenStudy (anonymous):

Yes 23x^2+5x-6732=0

OpenStudy (misty1212):

numbers too big for me i just did this http://www.wolframalpha.com/input/?i=x%2823x%2B5%29%3D6732

OpenStudy (anonymous):

When I did the quadratic formula I got 34

OpenStudy (misty1212):

you see you get \(17\)

OpenStudy (misty1212):

which is for some reason half of 17

OpenStudy (misty1212):

i mean half of your answer

OpenStudy (anonymous):

Hmmmm?

OpenStudy (misty1212):

did you check the link?

OpenStudy (anonymous):

Yes I did

OpenStudy (misty1212):

numbers too big for me to be interested in working out i just cheated one number is 17, the other is \(23\times 17+5=396\)

OpenStudy (anonymous):

I don't understand how I got 34! Lol

OpenStudy (misty1212):

the check is \(17\times 396=6732\) as needed

OpenStudy (misty1212):

maybe you divided incorrectly denominator of quadratic formula is \(2a\) i bet you divided by 23 not 46

OpenStudy (anonymous):

YEP! That is what I did!

OpenStudy (misty1212):

lol